# Intrinsic Current Concentration of Buffer Layer Material for Cable Ablation Failure: Role of Random Fiber Networks

### v1

@@ III Simulation Model
As depicted in Fig. 5 of a simulation model of $\rho=0.10$, the geometry scale of the whole model is $1000\,\upmu\mathrm{m}\times 500\,\upmu\mathrm{m}$. The fiber length is set as $70\,\upmu\mathrm{m}$ and the aspect ratio of each fiber is set as $70/3$. The conductivity of fiber is $10^{-2}$ S/m and the relative permittivity is set as 4741 [12]. Two rectangular electrodes are adjacent to the fiber networks, where the top electrode is the current input terminal with $1$ mA constant excitation and the bottom electrode is grounded. It should be noted that our simulation models were simplified, as the fiber size and the current excitation have been reduced. The reason for only applying 2D model is to save the calculation time.
@@ IV-A Current Distribution in the Buffer Layer
In the test platform shown in Fig. 3, the external current was applied to buffer layer under different compression states. The compression ratios of samples were respectively set to $10\%$ and $30\%$. And the surface temperatures were automatically recorded by the infrared camera in a period of time. Since the current distribution in buffer layer cannot be directly observed, it was thus reflected by the surface temperature distribution of the ITO glass. Fig. 6 shows the surface temperature distribution after 5 seconds of power-on, where the external current $I$ was set to $350$ mA. Significantly, a series of hot spots were captured by the infrared camera on surface. The hot spots represent the regions with higher temperatures. Fig. 6 was shot from directly above the test platform. From the top view, the current density in the buffer layer sample was perpendicular to the paper surface. Therefore, the current density of the hot spots regions must be higher than those of other regions. This reflects an important phenomenon, that is, after the buffer layer withstanding an external current, the current distribution inside the buffer layer is uneven and will be concentrated at certain sites. In the paper, we name this phenomenon as the intrinsic current concentration phenomenon. The so-called intrinsic refers that this phenomenon is caused by the buffer layer itself. And our paper thinks that the phenomenon is probably related to the microstructures of the random fiber networks of buffer layer.
@@ IV-C Current Distribution in the 2D Simulation Model
In order to investigate the relationships between the intrinsic current concentration and the random fiber networks, the current distribution inner the 2D simulation model, which was constructed based on the Mikado model, was calculated in this part. Fig. 9 shows the results of different fiber density. All the calculations has the same applied current excitation. It can be seen that the current distribution in the random fiber networks is uneven, where some fibers will withstand more current density than others. As presented in Fig. 9(a), the degree of current unevenness is more severe for lower density networks. In this case, the current will be concentrated on several fibers or the cross-linked regions of the fibers. It is conceivable that these sites will be priorly destroyed under the external current and they are the weaknesses of the fiber networks. Therefore, the simulation results can support the $\mu$CT results of weaknesses shown in Fig. 8.
@@
[…]
@@ Acknowledgments
The authors thank Ms. Guo and Mr. Ren at Instrument Analysis Center of Xi’an Jiaotong University for their assistance with LSCM and $\mu$CT analysis. The authors also thank the colleagues H. Sui and B. Li for their assistance with experiments.

### v2

@@ III Simulation Model
As depicted in Fig. 5 of a simulation model of $\rho=0.10$, the geometry scale of the whole model is 1000 $\upmu$m×$\times$500 $\upmu$m. The fiber length is set as 70 $\upmu$m and the aspect ratio of each fiber is set as 70/3. The conductivity of fiber is 10-2 S/m and the relative permittivity is set as 4741 [12]. Two rectangular electrodes are adjacent to the fiber networks, where the top electrode is the current input terminal with 1 mA constant excitation and the bottom electrode is grounded. It should be noted that our simulation models were simplified, as the fiber size and the current excitation have been reduced. The reason for only applying 2D model is to save the calculation time.
@@ IV-A Current Distribution in the Buffer Layer
In the test platform shown in Fig. 3, the external current was applied to buffer layer under different compression states. The compression ratios of samples were respectively set to 10% and 30%. And the surface temperatures were automatically recorded by the infrared camera in a period of time. Since the current distribution in buffer layer cannot be directly observed, it was thus reflected by the surface temperature distribution of the ITO glass. Fig. 6 shows the surface temperature distribution after 5 seconds of power-on, where the external current $I$ was set to 350 mA. Significantly, a series of hot spots were captured by the infrared camera on surface. The hot spots represent the regions with higher temperatures. Fig. 6 was shot from directly above the test platform. From the top view, the current density in the buffer layer sample was perpendicular to the paper surface. Therefore, the current density of the hot spots regions must be higher than those of other regions. This reflects an important phenomenon, that is, after the buffer layer withstanding an external current, the current distribution inside the buffer layer is uneven and will be concentrated at certain sites. In the paper, we name this phenomenon as the intrinsic current concentration phenomenon. The so-called intrinsic refers that this phenomenon is caused by the buffer layer itself. And our paper thinks that the phenomenon is probably related to the microstructures of the random fiber networks of buffer layer.
@@ IV-C Current Distribution in the 2D Simulation Model
In order to investigate the relationships between the intrinsic current concentration and the random fiber networks, the current distribution inner the 2D simulation model, which was constructed based on the Mikado model, was calculated in this part. Fig. 9 shows the results of different fiber density. All the calculations has the same applied current excitation. It can be seen that the current distribution in the random fiber networks is uneven, where some fibers will withstand more current density than others. As presented in Fig. 9(a), the degree of current unevenness is more severe for lower density networks. In this case, the current will be concentrated on several fibers or the cross-linked regions of the fibers. It is conceivable that these sites will be priorly destroyed under the external current and they are the weaknesses of the fiber networks. Therefore, the simulation results can support the $\upmu$CT results of weaknesses shown in Fig. 8.
@@
$I_{i}=\int_{l_{i}}\,{\textbf{{J}}}\cdot{\textbf{{n}}}\,\mathrm{d}l\quad(i=1,2,\,\cdots,10)$ (5)
@@ Acknowledgments
The authors thank Ms. Guo and Mr. Ren at Instrument Analysis Center of Xi’an Jiaotong University for their assistance with LSCM and $\upmu$CT analysis. The authors also thank the colleagues H. Sui and B. Li for their assistance with experiments.